# Vibe coding: programming through conversation with artificial intelligence

### v1

@@ Abstract
We examine “vibe coding,” an emergent programming paradigm where developers primarily write code by interacting with code-generating large language models rather than writing code directly. We analysed a curated set of videos depicting extended vibe coding sessions with rich think-aloud reflections. Using framework analysis, we investigated programmers’ goals, workflows, prompting techniques, debugging approaches, and challenges encountered.
We find that vibe coding follows iterative goal satisfaction cycles where developers alternate between prompting AI, evaluating generated code through rapid scanning and application testing, and manual editing. Prompting strategies blend vague, high-level directives with detailed technical specifications. Debugging remains a hybrid process combining AI assistance with manual practices.
Critically, vibe coding does not eliminate the need for programming expertise but rather redistributes it toward context management, rapid code evaluation, and decisions about when to transition between AI-driven and manual manipulation of code. Trust in AI tools during vibe coding is dynamic and contextual, developed through iterative verification rather than blanket acceptance. Vibe coding is an evolution of AI-assisted programming that represents an early manifestation of “material disengagement”, where practitioners orchestrate code production and manipulation, mediated through AI, while maintaining selective and strategic oversight.
@@ 1 Introduction and Background
This description of a new style of programming, that shall henceforth be referred to as the Karpathy canon, has had significant influence. Beyond sparking discourse, it has spurred the adoption of the term “vibe coding” to describe an idealised style of programming involving the use of agentic code-generating language model tools such as Cursor, GitHub Copilot Agent mode, Windsurf, Bolt, and others. From the excerpt above, it would seem that the core ideal of vibe coding is disengagement from directly working with code (authoring, editing, reading, etc.) and instead trusting the agentic tool to perform these operations reliably enough for practical use, based on natural language (preferably dictated) descriptions of the vibe coder’s intent.
@@
What often occurs following such philosophical seeding is the collective mobilisation of the community of programmers to make sense of, adapt, and adopt (or resist) this proposed way of working. For some, the Karpathy canon merely provides a convenient label for a practice they were already engaged in. For others, the term is an introduction to a new way of working, but with innumerable unanswered questions that need to be negotiated through personal practice on a case-by-case basis. Is it still vibe coding if I manually edit the code from time to time? Is it still vibe coding if I use the keyboard instead of dictation? Is it vibe coding if I use a non-agentic tool? What is the spirit of vibe coding, and how does it differ from its letter?
The manner by which a community, through evolving practice and discourse, comes to hash out these questions and elaborate the concept of vibe coding is a sociological question, which we will not attempt to address in this paper. We raise this to emphasise that, despite nucleating around the Karpathy canon, vibe coding is an emergent phenomenon and its properties are being continuously negotiated as of this writing.
@@
Thus motivated, we contribute the first empirical analysis of vibe coding as an emerging programming phenomenon, through a framework analysis of curated think-aloud videos of programmers vibe coding sourced from YouTube and Twitch (Section 2). Our analysis provides a preliminary picture of how goals and intentions are formed during vibe coding, the vibe coding workflow, the nature of prompting, debugging practices, challenges that arise during vibe coding, what kinds of programmer expertise are deployed and how, and the nature of trust in AI during vibe coding (Section 3). We discuss how vibe coding appears to represent an evolution of previous generations of AI-assisted coding, the vibe “gestalt”, the tradeoffs of material disengagement, and how vibe coders attempt to influence societal attitudes to the use of AI (Section 4).
Our analysis is preliminary and provisional; we study only a small set of videos in detail, and the landscape of tools and practices associated with vibe coding is rapidly evolving. Our findings should therefore be understood as a temporally contingent snapshot of how vibe coding appears to manifest at this specific early juncture. Such a vignette serves both as a foundation for future work on understanding the vibe coding phenomenon, and as a reference point for comparison as the practice evolves further.
@@ 2.2 Video analysis method
Given that vibe coding is an emergent phenomenon, we sought the flexibility to document isolated events, sequential event chains, concurrent events, variations in the duration of phenomena of interest, and temporally disconnected but conceptually related events (e.g., when a programmer initiates an activity, transitions to another task, and subsequently resumes the original work). Thus we chose not to segment videos into predefined episodes or fixed time intervals (e.g., 20-second segments, as exemplified in Srinivasa Ragavan et al., (2021)). Moreover, we required a methodology that permitted flexible analysis of both verbal think-aloud information and observed actions within the videos.
@@
Unlike codes in a qualitative codebook, framework categories do not represent concrete phenomena but rather constitute conceptual domains into which observations can be classified. This approach thus requires less initial commitment to an ontology of findings. Rather, the indexing process serves to reorganise the data (they are used merely to collate data about a concept of interest, rather than data in support of a phenomenal theme) so that the directed research questions implied by the framework categories can then be addressed through various interpretive methods. In our case, because the size of recorded observations in each category was relatively small, two researchers were able to collaboratively negotiate the most salient findings from each category.
@@ 2.3 Video selection
We selected an initial longlist of videos by searching the YouTube and Twitch platforms using platform native search using the search terms “vibe coding” and “vibe coding session”. We considered but refrained from including a wider array of search terms (e.g., including tool names such as “Cursor”) because as mentioned above, vibe coding is an emergent practice that is being continuously defined and negotiated and our study is partly an exercise in understanding that process; we thus adopt the methodological principle of ordinary language philosophy (Laugier,, 2013) of attending to the term as it is used and enacted in practice, rather than attempting to impose any a priori interpretations of vibe coding, as including other keywords would have done.
Searches were conducted during the period 17 March 2025 and 02 April 2025. Prior to inclusion, videos were lightly inspected to exclude any false positive videos that clearly contained no developer reflections about the concept of vibe coding (e.g., advertisements, slideshows with text-to-speech voiceovers, ambient music videos, etc.) and extremely short videos (less than 3 minutes). This resulted in an initial list of 35 videos with a combined duration of 35 hours, 46 minutes and 21 seconds.
Next, the two researchers independently viewed 10-20 minute sections of each video to sensitise and familarise themselves with the content, making initial notes about potential themes and quality criteria. Through negotiation, four videos were selected for the two researchers to view in their entirety to perform a preliminary qualitative analysis.
@@ 2.4 Framework analysis
For determining the categories of the framework analysis, the researchers independently compiled notes while viewing videos from the entire longlist on potential categories and research questions to address, resulting in a combined list of 22 proto-categories at various levels of granularity. After negotiation, three categories (on programmer goals, the formation of intention, and the deployment of expertise) were further elaborated to accommodate nuances that researchers had observed at this initial stage. Two proto-categories were merged into a broader category on workflows and tools, and one proto-category was removed, namely, how programmers and non-programmers differ in their approach to vibe coding, because our dataset did not contain any vibe coding sessions featuring a user with no programming expertise. Our final framework consisted of 9 top-level categories organised into 20 subcategories. This full framework is given in Table LABEL:tab:framework.
@@
Finally, the two researchers divided the dataset into two parts of similar duration, and independently applied the framework to qualitatively analyse one part each. In a slight deviation from the standard method of framework analysis where units of the analysed data are directly attached (“indexed”) to framework categories, in this case the researchers made qualitative notes and observations, including excerpts from the think aloud transcripts of the videos, under each category as appropriate. This deviation is necessary to accommodate the difficulties with unitising videos to properly account for emergent phenomena (as mentioned earlier) and also to account for the fact that a phenomenon may be observed by the researcher as manifesting through an irreducible combination of onscreen actions and think-aloud reflection (which may not occur contemporaneously).
The results of this analysis are presented in Section 3.
Top Level Category Subcategories Elaboration
Goals a. What do they build? i. What kind of apps do people build using vibe coding?
Intentions a. Initial intention? i. How is the intention for an app formed and refined in vibe coding?
b. Intention refined? How and why? i. What AI workflow resources (e.g., exploration, seeking alternatives) support and refine this intent?
Workflow a. Stages of workflow? i. What are the stages of the vibe coding workflow?
b. Portfolio of tools and technologies? i. How do different tools work together in the context of vibe coding?
b. Explicit prompting strategies? i. What are the prompting strategies in vibe coding?
c. Granularity – high/low level? i. What is the granularity of prompts and iterations?
d. Single/multi-objective prompts? i. Do users address multiple issues at once or handle one problem at a time?
e. Modes – voice/text/images? i. What is the role of different modes (voice, text, screenshot, etc.) in vibe coding?
Debugging a. What debugging strategies for their target program? i. What debugging strategies are employed in vibe coding?
Challenges a. What challenges (besides code bugs) in achieving goals? i. What challenges and barriers are encountered during vibe coding?
b. How addressed? i. What strategies are adopted to address these challenges?
ii. In what ways do experts deploy their expertise?
b. What kinds of expertise used? i. Is the relevant expertise strictly developer knowledge, or does it include understanding system requirements, end-to-end design, and the broader technical ecosystem?
c. When switch to manual work? i. When do practitioners transition from vibe coding to manual editing (code, text, image, artifact)?
Definition of vibe coding and performance a. How programmers define vibe coding? i. When is an activity considered vibe coding versus simply using AI tools?
b. How performance for YouTube/Twitch skews production? i. How does the performative aspect of vibe coding (e.g., YouTube demonstrations) influence the practice and perception of generative AI in programming?
@@ 3.1.1 What kinds of applications are built using vibe coding?
Our examples spanned varying domains and complexity levels. In the simplest case (YT15), a creator developed an animated diagram explaining a coding agent’s workflow. This project focused primarily on visual design elements, including arrow positioning, labelling, animation, and aesthetic styling with specific colour themes. In TW1, the creator created and successfully deployed various member-exclusive features to an existing gaming website.
The YT21 creator built a code-base chat application that integrated with GitHub repositories, featuring user authentication, persistent chat histories across sessions, and URL-based navigation to specific conversations. The project applied retrieval augmented generation to answer questions about a user’s codebase. In YT22, a developer created a web application to help international students with job applications, incorporating resume tailoring based on job descriptions, user account management, payment processing via Stripe, and file upload functionality.
@@
TW1 also expected full success for vibe coding new features to add to her deployed website. Rather than leveraging vibe coding in an exploratory manner, her approach combined vibe coding to generate code for each feature with manual fixes and programming to accomplish a set of tasks she had collected on a to-do list.
@@ 3.2.2 How and why is intention refined over the course of the vibe coding session?
The iterative, conversational nature of vibe coding supports the refinement of intention through progressive discovery and refinement. Across all videos, the creator’s ideas expanded beyond their initial vision as refinements were driven by factors such as dissatisfaction with the initial output and emerging needs or ideas during the process of evaluating the output. In the vibe coding workflow, users must simultaneously form intentions for the artifact they are developing and form intentions for how best to communicate those goals to the AI effectively. The inevitable mismatch between the user’s internal intent and the AI’s interpretation of it becomes a significant driver for refining not only the project’s direction but also the user’s prompting strategies and their understanding of the AI’s capabilities.
AI contributes to shaping final intent by filling implementation gaps and suggesting new possibilities. In YT15, the intent clearly evolved through multiple iterations, starting with an initial diagram layout, developing the placement of arrows, then adding styling and animation. The direction of the project was significantly shaped by the back-and-forth with the AI as the generated code variously: succeeded in matching the creator’s intent directly, thus allowing progress; or succeeded indirectly such as by interpreting vague instructions, thus sparking ideas; or failed, thus necessitating debugging or a goal “pivot” (Drosos et al.,, 2024).
In one instance for TW1, the AI generated code had an error, but on inspection she realised there was a bug in the previous code that was revealed by the new code. In another instance, the code generated did not match her original intent, but gave her ideas for how it might be useful for implementing future alternatives.
@@
AI is also consulted outside of the vibe coding process for guidance and iteration on intent, such as soliciting feature ideas (TW1). Other traditional sources of intention refinement, such as viewing prior art and alternative commercial products as design references, was also observed (YT21).
The blending of prototyping and production, and context momentum
@@ 3.3.1 What are the stages of the vibe coding workflow?
The vibe coding workflow consistently begins with an initial goal-setting and prompting phase. In YT15, this involved the creator giving a detailed spoken prompt to Cursor via a transcription tool to generate the initial diagram, establishing a diagram explaining the code agent’s workflow as the first goal. Similarly, the creator in YT21 started with a clear goal: to modify his existing ‘chat with a code base’ application to support multiple chat conversations per repository, which they subsequently decomposed into smaller sub-goals. In YT22, the creator sketched a flow diagram to visualise the tool’s functionality and applied Code Guide, [^2] https://www.codeguide.dev/ a tool that generated a project plan and documentation from his app requirements.
@@
Thus, vibe coding is characterised by iterative goal satisfaction (Drosos et al.,, 2024) cycles, where developers (1) Formulate a goal or sub-goal, (2) Prompt the model to generate code to achieve that goal, (3) Review the generated code, (4) Accept or reject the changes, (5) Test the changes in the application, (6) Identify any bugs or areas for improvement, (7) Refine the prompt or transition to manual debugging and editing, (8) Repeat until the sub-goal was satisfied, refined, or pivoted.

### v2

@@ Abstract
We examine “vibe coding”: an emerging programming paradigm where developers primarily write code by interacting with code-generating large language models rather than writing code directly. We present the first empirical study of vibe coding. We analysed over 8 hours of curated video capturing extended vibe coding sessions with rich think-aloud reflections. Using framework analysis, we investigated programmers’ goals, workflows, prompting techniques, debugging approaches, and challenges encountered.
We find that vibe coding follows iterative goal satisfaction cycles where developers alternate between prompting AI, evaluating generated code through rapid scanning and application testing, and manual editing. Prompts in vibe coding blend vague, high-level directives with detailed technical specifications. Debugging remains a hybrid process combining AI assistance with manual practices.
Critically, vibe coding does not eliminate the need for programming expertise but rather redistributes it toward context management, rapid code evaluation, and decisions about when to transition between AI-driven and manual manipulation of code. Trust in AI tools during vibe coding is dynamic and contextual, developed through iterative verification rather than blanket acceptance. Vibe coding is an evolution of AI-assisted programming that represents an early manifestation of “material disengagement”, wherein practitioners orchestrate code production and manipulation, mediated through AI, while maintaining selective and strategic oversight.
Category Associated questions
Goals What kinds of applications are vibe coded? What degree of success is expected? To what extent is exploration a motivation? (Section 3.1)
Intentions How are implementation ideas initially formed and refined? Is intent pre-defined or shaped iteratively through interaction with AI? (Section 3.2)
Workflow What are the stages of the vibe coding workflow? How is programmer effort distributed? How are various tools and resources combined? (Section 3.3)
Prompting How do programmers express their intent to the system? What setup and prompting strategies are used? What is the granularity of instructions? Are prompts single- or multi-objective? What input modes are employed? (Section 3.4)
Debugging What methods are used to identify and resolve problems in AI-generated code? (Section 3.5)
Challenges What technical, conceptual, and workflow-related obstacles arise in vibe coding beyond debugging? What strategies are used to overcome them? (Section 3.6)
Expertise What forms of knowledge are deployed in vibe coding? When do practitioners transition from instructing AI to direct manual work? (Section 3.7)
Trust How do users build confidence in AI outputs? What are the risks of overreliance? (Section 3.8)
Definition and Performance How is vibe coding understood relative to other AI-assisted programming practices? How do performative contexts such as video streaming shape the practice and its perception? (Section 3.9)
Table 1: Major categories of findings from our study of the practice of vibe coding. Findings were derived from in-depth qualitative framework analysis of approximately 8.5 hours of video of extended think-aloud vibe coding sessions.
@@ 1 Introduction and Background
This description of a new style of programming, that shall henceforth be referred to as the Karpathy canon, has had significant influence. Beyond sparking discourse, it has spurred the adoption of the term “vibe coding” to describe an idealised style of programming involving the use of “agentic” [^1] Unfortunately, there is no consensus on the definition of the term “agentic” or “agent” in this context. It is primarily a commercial term rather than a technical term. We invoke this term to make an explicit connection with that commercial discourse, and do not make ontological claims about “agents” as a clearly distinct category of AI tools. For our purposes, the property of interest is that tools described as “agentic” can typically take a greater range of programmatic actions on the user’s content (such as reading code, making direct edits across multiple files, and executing terminal commands) than previous implementations of AI assistance within software development environments. software development tools such as Cursor, GitHub Copilot Agent mode, Windsurf, Bolt, and others. From the excerpt above, it would seem that the core ideal of vibe coding is disengagement from directly working with code (authoring, editing, reading, etc.) and instead trusting the agentic tool to perform these operations reliably enough for practical use, based on natural language (preferably dictated) descriptions of the vibe coder’s intent.
@@
What often occurs following such philosophical seeding is the collective mobilisation of the community of programmers to make sense of, adapt, and adopt (or resist) this proposed way of working. For some, the Karpathy canon merely provides a convenient label for a practice they were already engaged in. For others, the term is an introduction to a new way of working, but with innumerable unanswered questions that need to be negotiated through personal practice on a case-by-case basis. Is it still vibe coding if I manually edit the code from time to time? Is it still vibe coding if I use the keyboard instead of dictation? Is it vibe coding if I use a non-agentic tool? How do I vibe code effectively? What is the spirit of vibe coding, and how does it differ from its letter?
The manner by which a community, through reflexive practice and discourse, comes to hash out these questions and elaborate the concept of vibe coding is a sociological question, which we will not attempt to address in this paper. We raise this to emphasise that, despite nucleating around the Karpathy canon, vibe coding is an emerging phenomenon and its properties are being continuously negotiated as of this writing.
@@
Thus motivated, we contribute the first empirical analysis of vibe coding as an emerging programming phenomenon, through a framework analysis of curated think-aloud videos of programmers vibe coding sourced from YouTube and Twitch (Section 2). Our analysis provides a picture of how goals and intentions are formed during vibe coding, the vibe coding workflow, the nature of prompting, debugging practices, challenges that arise during vibe coding, what kinds of programmer expertise are deployed and how, and the nature of trust in AI during vibe coding (Section 3). We discuss how vibe coding appears to represent an evolution of previous generations of AI-assisted coding, the vibe “gestalt”, the tradeoffs of material disengagement, and how vibe coders attempt to influence societal attitudes to the use of AI (Section 4).
Our analysis is preliminary; we study only a small set of videos in detail, and the landscape of tools and practices associated with vibe coding is rapidly evolving. Our findings should therefore be understood as a temporally contingent snapshot of how vibe coding appears to manifest at this specific early juncture. Such a vignette serves both as a foundation for future work on understanding the vibe coding phenomenon, and as a reference point for comparison as the practice evolves further.
@@ 2 Method
Figure 1: Overview of research process. An initial longlist of 35 videos was analysed in two phases, resulting in a framework for qualitative analysis and a set of inclusion criteria. Applying these criteria, 5 videos (approximately 8.5 hours) were analysed in detail in their entirety according to the framework.
@@ 2.2 Video analysis method
Given that vibe coding is an emerging phenomenon, we sought the flexibility to document isolated events, sequential event chains, concurrent events, variations in the duration of phenomena of interest, and temporally disconnected but conceptually related events (e.g., when a programmer initiates an activity, transitions to another task, and subsequently resumes the original work). Thus we chose not to segment videos into predefined episodes or fixed time intervals (e.g., 20-second segments, as exemplified in Srinivasa Ragavan et al., (2021)). Moreover, we required a method that permitted flexible analysis of both verbal think-aloud information and observed actions within the videos.
@@
Unlike codes in a qualitative codebook, framework categories do not represent concrete phenomena but rather constitute conceptual domains into which observations can be classified. This approach thus requires less initial commitment to an ontology of findings. Rather, the indexing process serves to reorganise the data (they are used merely to collate data about a concept of interest, rather than data in support of a phenomenal theme) so that the directed research questions implied by the framework categories can then be addressed through various interpretive methods. In our case, because the size of recorded observations in each category was relatively small (around 2000 words of combined researcher notes per category), two researchers were able to collaboratively negotiate the most salient findings from each category.
@@ 2.3 Video selection
We selected an initial longlist of videos by searching the YouTube and Twitch platforms using platform native search using the search terms “vibe coding” and “vibe coding session”. We considered but refrained from including a wider array of search terms (e.g., including tool names such as “Cursor”) because as mentioned in Section 1, vibe coding is an emerging practice that is being continuously defined and negotiated and our study is partly an exercise in understanding that process; we thus adopt the methodological principle of ordinary language philosophy (Laugier,, 2013) of attending to the term as it is used and enacted in practice, rather than attempting to impose any a priori interpretations of vibe coding, as including other keywords would have done.
Searches were conducted during the period 17 March 2025 to 02 April 2025. Prior to inclusion, videos were lightly inspected to exclude any false positive videos that clearly contained no developer reflections about the concept of vibe coding (e.g., advertisements, slideshows with text-to-speech voiceovers, ambient music videos, etc.) and extremely short videos (less than 3 minutes). This resulted in an initial list of 35 videos with a combined duration of 35 hours, 46 minutes and 21 seconds.
Next, two researchers independently viewed 10-20 minute sections of each video to sensitise and familarise themselves with the content, making initial notes about potential themes and quality criteria. Through negotiation, four videos were selected for the two researchers to view in their entirety to perform a preliminary qualitative analysis.
@@ 2.4 Framework analysis
For determining the categories of the framework analysis, the researchers independently compiled notes while viewing videos from the entire longlist on potential categories and research questions to address, resulting in a combined list of 22 proto-categories at various levels of granularity. After negotiation, three categories (on programmer goals, the formation of intention, and the deployment of expertise) were further elaborated to accommodate nuances that researchers had observed at this initial stage. Two proto-categories were merged into a broader category on workflows and tools, and one proto-category was removed, namely, how programmers and non-programmers differ in their approach to vibe coding, because our dataset did not contain any vibe coding sessions featuring a user with no programming expertise.
Our final framework consisted of 9 top-level categories organised into 20 subcategories. The 9 top-level framework categories were: goals, intentions, workflow, prompting, debugging, challenges, expertise, trust, and definition and performance. An overview of these categories and some of the questions associated with each is given in Table 1. The full framework, including subcategories and elaborations of specific questions, is given in Appendix B.
@@
Finally, two researchers divided the dataset into two parts of similar duration, and independently applied the framework to qualitatively analyse one part each. In a slight deviation from the standard method of framework analysis where units of the analysed data are directly attached (“indexed”) to framework categories, in this case the researchers made qualitative notes and observations, including excerpts from the think aloud transcripts of the videos, under each category as appropriate. This deviation is necessary to accommodate the difficulties with unitising videos to properly account for emerging phenomena (as mentioned earlier) and also to account for the fact that a phenomenon may be observed by the researcher as manifesting through an irreducible combination of onscreen actions and think-aloud reflection (which may not occur simultaneously). An overview of the research process is given in Figure 1. The results of this analysis are presented in Section 3.
@@ 3.1.1 What kinds of applications are built using vibe coding?
Our examples spanned varying domains and complexity levels. In the simplest case (YT15), a creator developed an animated diagram explaining a coding agent’s workflow. This project focused primarily on visual design elements, including arrow positioning, labelling, animation, and aesthetic styling with specific colour themes. In TW1, the creator created and successfully deployed various member-exclusive features to an existing gaming website. The YT21 creator built a code-base chat application that integrated with GitHub repositories, featuring user authentication, persistent chat histories across sessions, and URL-based navigation to specific conversations. The project applied retrieval augmented generation to answer questions about a user’s codebase. In YT22, a developer created a web application to help international students with job applications, incorporating resume tailoring based on job descriptions, user account management, payment processing via Stripe, and file upload functionality.
@@
TW1 also expected full success for vibe coding new features to add to their deployed website. Rather than applying vibe coding in an exploratory manner, their approach combined vibe coding to generate code for each feature with manual fixes and programming to accomplish a set of tasks they had collected on a to-do list.
@@ 3.2.2 How and why is intention refined over the course of the vibe coding session?
The iterative, conversational nature of vibe coding supports the refinement of intention through progressive discovery and refinement. Across all videos, the creator’s ideas expanded beyond their initial vision as refinements were driven by factors such as dissatisfaction with the initial output and emerging needs or ideas during the process of evaluating the output. In the vibe coding workflow, users must simultaneously form intentions for the artefact they are developing and form intentions for how to communicate those goals effectively. The inevitable mismatch between the user’s internal intent and the AI “interpretation” of it becomes a significant driver for refining not only the project’s direction but also the user’s prompting strategies and their understanding of AI capabilities.
AI contributes to shaping final by filling implementation gaps and suggesting new possibilities. In YT15, the intent clearly evolved through multiple iterations, starting with an initial diagram layout, developing the placement of arrows, then adding styling and animation. The direction of the project was significantly shaped by the back-and-forth with the system as the generated code variously: succeeded in matching the creator’s intent directly, thus allowing progress; or succeeded indirectly such as by interpreting vague instructions, thus sparking ideas; or failed, thus necessitating debugging or a goal “pivot” (Drosos et al.,, 2024).
In one instance for TW1, the AI generated code had an error, but on inspection they realised that there was a bug in the previous code that was revealed by the new code. In another instance, the code generated did not match their original intent, but gave them ideas for how it might be useful for implementing future alternatives.
@@
AI tools are also consulted outside of the vibe coding process for guidance and iteration on intent, such as soliciting feature ideas (TW1). Other traditional sources of intention refinement, such as viewing prior art and alternative commercial products as design references, were also observed (YT21).
The blurring of prototyping and production.
@@ 3.3.1 What are the stages of the vibe coding workflow?
The vibe coding workflow consistently begins with an initial goal-setting and prompting phase. In YT15, this involved the creator giving a detailed spoken prompt to Cursor via a transcription tool to generate the initial diagram, establishing a diagram explaining the code agent’s workflow as the first goal. Similarly, the creator in YT21 started with a clear goal: to modify his existing ‘chat with a code base’ application to support multiple chat conversations per repository, which they subsequently decomposed into smaller sub-goals. In YT22, the creator sketched a flow diagram to visualise the tool’s functionality and applied Code Guide, [^3] Available at: https://www.codeguide.dev/. Last accessed: 19 September 2025. a tool that generated a project plan and documentation from his app requirements.
@@
Thus, vibe coding is characterised by iterative goal satisfaction (Drosos et al.,, 2024) cycles, where developers (1) Formulate a goal or sub-goal, (2) Prompt the model to generate code to achieve that goal, (3) Review the generated code, (4) Accept or reject the changes, (5) Test the changes in the application, (6) Identify any bugs or areas for improvement, (7) Refine the prompt or transition to manual debugging and editing. Steps (2)-(7) are repeated until the sub-goal is satisfied or pivoted, at which point the programmer returns to step (1). The entire process is repeated until the programmer decides not to formulate any further goals.